Services Unit Meeting 06/10/2009

Present: Craig, Neil, Gordon, Roger
Room: IF 4.02

Sleekit AFS web

The temp fix is in place, but it hight lighted a problem in the apache component, which has been fixed by hand on sleekit for now, but needs a new apache component to be shipped.

Cosign V3

groups.inf.ed.ac.uk  - done
ifile.inf.ed.ac.uk   - done
jabber.inf.ed.ac.uk  - done
jings.inf.ed.ac.uk   - done
lists.inf.ed.ac.uk   - done
rbs.inf.ed.ac.uk
rbstest.inf.ed.ac.uk
wcms.inf.ed.ac.uk    - done

Should manage to meet tomorrows deadline for all but the rbs stuff, we'll ask for some leeway.

Robusitication

mullo - a date has not been agreed for a reboot. We'll suggest that we just do it in the small hours.

still need to multipath/network bond various file servers

Server Room moves

Sun said we can move it, but we'll be billed from any damage caused during the move. KB move probably in the next 2 weeks.

BP - we believe hippocampus and pegasus are clear of data. So only the macbackup retrospect service to move, and the BP web server, sleekit, service needs to be moved. Mirrors need to be moved off pegasus/hippo.

FC5 machines

maelcum, getting close to actually looking at this.

powell, linotype - not managed to speak to the efiancials people. Also need network holes for port 9100 on the printers and an entry to route traffic to the printer wire from efiancials.

Buying printers

Alastair has been off, so Craig's not spoken to him. Gordon's done some quick sums, and reckons the new printers are working out at worst about 2p side.

Printing from wireless

Not added VPN to CUPs wires yet. Some concern about anonymity of print jobs from selfmanaged machines.

Moving server configs

Doing it when we happen to be looking at machines/services for other reasons.

Roger is still to look at wiki.

twiki to confluence

Roger installed a test confluence on his desktop and tried one of the tools to convert the DICE TWiki into confluence, it just worked for (very ball park figure) ~70% of the content. It could take several person days to review what was converted and fix up broken things (again a very ball park figure).

New Web Server hardware

Again Alastair was away.

Pandemic

Neil is suggested that some how people indicate that they are happy with their pandemic "training". He's been marking himself in bold on the Pandemic page.

George needs to be added to the networker mailing list.

Keys - need to make sure if keys/access is required to get to machines, that people know how to get that access/keys. eg swipe card access to BP. FH corridor access.

dandy smtp.inf

Hasn't been relocated to the new rack, but will be robustified.

cameleopard memory

Neil has been chasing Dell, but they are not getting back to him.

Timekeeping

Craig still to get Q3 figures.

Have not had buckets meeting.

www.cogsci

No progress on this.

AFS DB hardware

Still waiting for the go-ahead as where to put them. Need to move BP AFS DB to forum. Perhaps we should just resurrect some old 1U machines eg 750s.

Mermaid

Still not done.

-- CraigStrachan - 06 Oct 2009

Topic revision: r5 - 04 Nov 2009 - 15:19:27 - RogerBurroughes
 
This site is powered by the TWiki collaboration platformCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback
This Wiki uses Cookies